OpenStudy (anonymous):

The ellipse with x-intercepts (5, 0) and (-5, 0); y-intercepts (0, 3) and (0, -3).

OpenStudy (anonymous):

g'day....:) what u want... the equation?

OpenStudy (anonymous):

you'll have to identify a, b then put it in this form: \[\huge \frac{x^2}{a^2}+\frac{y^2}{b^2}=1 \] where 2a is the length of the major axis, and 2b is the length of the minor axis

OpenStudy (anonymous):

if y=0, then \[ \frac {x^2}{a^2}=1\\ x=\pm a \]

OpenStudy (anonymous):

If x =0, then \[y=\pm b \]

OpenStudy (anonymous):

What is a and b?

OpenStudy (anonymous):

Thats exactly what Im having a problem with...finding a,b

OpenStudy (anonymous):

Did you see my posts above?

OpenStudy (anonymous):

How much is y in (5, 0) and (-5, 0);

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

i think it would be positive negative 25 for y

OpenStudy (anonymous):

a=5 and b=3

OpenStudy (anonymous):

\[ \frac {x^2}{25} + \frac {y^2}9=1 \]

OpenStudy (anonymous):

yes so thats the equation I need right

OpenStudy (anonymous):

Thank you sooo much...you are very patient